Two boxes are connected by a light string that passes over a light, frictionless pulley. One that rises at 30.0 degrees above the horizontal, and the system isreleased from rest. Find the acceleration of each box? 50.kg box and 30.0kg box
Answer #1
Given that the masses are M1 = 50.0 Kg and M2 = 30.0 KgThe angle of inclination is θ = 30.00------------------------------------------------------------------------------------Let the mass M1 = 50.0 Kg is moving upwards thenthe situation is as shown in the below figure.since the bodies are moving with the accelaration a then apply Newtons law for eachbody thenFor the mass M1:T-M1gsinθ = M1*a --------------(1)For the mass M2:M2g- T = M2*aT=M2g -M2*a--------------(2)From the equations (1) and (2) we getM2g-M1*a- M1gsinθ = M1*aM2g-M1gsinθ = M1*a+M2*athen the accelaration a =( M2g - M1gsinθ)/(M1 +M2)=(30*9.8-50*9.8*sin30)/80=0.6125m/s2Since a is positive, the mass M1 moving upwards and themassM2 is moving downwards